Graph the curve whose parametric equations are given. x = t3 + 1, y = t3 - 10;-2 sts2 ++++ 30 x -30 -10 10 x -10 O
Answer #1

Given x ² t tt and y = t3 to then T-X - et put 3 in ot-g7 y= OT-T-X = y = x - 1L ay ری Х (1,0) (ant